Circular No. 7756

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1039 CENTAURI = POSSIBLE NOVA IN CENTAURUS M. Della Valle, Arcetri Astrophysical Observatory, Florence; C. Maras-

ton, Munich Observatory; O. Hainaut, European Southern Observatory; and R. Williams, Space Telescope Science Institute, report: “Preliminary analysis of a spectrum (range 400–900 nm) of the suspected nova discovered by Liller in October (IAUC 7726, 7727, 7731), obtained on Nov. 12.4 UT at La Silla with the New Technology Telescope (+ EMMI), shows this object indeed to be a classical nova, currently in transition to the nebular phase of its decline. The spectrum is dominated by Hα, O I (mult. 34), Mg I (mult. 7), Mg II (mult. 8), Fe II (mult. 42, 48, 73), and [N II] and [O I] emission lines. The nova is therefore of the Fe II class, and the line width measured for Hα yields FWZI = 4000 km/s.”

N. N. Samus, Institute of Astronomy, Moscow, informs us that the designation V1039 Cen has been given to this nova.

NOVA IN M33 O. Trondal, Oslo, Norway, reports his discovery of an apparent nova

(mag ≈ 16.5) on a CCD image taken on Nov. 15.762 UT with a 0.25-m f/4.2 reflector. The new object is located at α = 1^{h}33^{m}42^{s}.84, δ = +30^{o}40′53′′.0 (equinox 2000.0). Nothing was present at this location on an image taken by Trondal on Nov. 13.92. W. Li, University of California at Berkeley, writes that the new object, located ∼ 104′′ west and 76′′ north of the nucleus of M33 (= NGC 598), is also confirmed (mag ∼ 17) on an image taken by M. Schwartz with the Tenagra Observatory 0.5-m telescope on Nov. 17.2; an image taken with the same telescope on Nov. 12.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001fv IN NGC 3512 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001fv (cf. IAUC 7750), obtained by P. Berlind on Nov. 14.52 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova well past maximum light. The spectrum shows a continuum with well-defined P-Cyg lines of hydrogen, sodium, and iron, consistent with the plateau phase of a normal type-II supernova. Adopting the NASA/IPAC Extragalactic Database recession velocity of 1376 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 3500 km/s.

2001 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green